### v2

@@ 5 Results and discussion
This is important for an accurate description of the short-range electron-nucleon interactions. However, in our case, because of the significant long-range contribution, this region contribution is negligible. This is in sharp contrast to the eN-SPS interaction.
The equilibrium geometry was determined in [30] to correspond to $R=4.4\,{\rm a.u.}$ The property was computed on a grid of values of $R$ and $\theta$. $R$ was taken to span values from $3.6\,{\rm a.u.}$ to $6.0\,{\rm a.u.}$ with step $0.2\,{\rm a.u.}$, where $\theta$ was taken to be $0^{\circ}$, $25^{\circ}$, $57^{\circ}$, $90^{\circ}$, $122^{\circ}$, $155^{\circ}$ and $180^{\circ}$. The resulting value for the equilibrium geometry equals $1.1407\cdot 10^{-5}\lambda_{e}^{-1}$ where $\lambda_{e}=\frac{\hbar}{m_{e}c}$ is the electron Compton length. Compare this with $3.36\cdot 10^{-5}\lambda_{e}^{-1}$ for YbOH obtained in [26]. The dependence on the $\theta$ and $R$ is depicted on the Fig. 3 and Fig. 4. The values for several rovibrational states are given in Table 1. To check the convergence and estimate numerical error the computation was performed for $l_{max},j_{max},k_{max}=40,\ldots 43$, and $n_{max}=10,\ldots 13$.
@@
The YbOH equilibrium configuration (FS-CCSD, $m_{a}=1..10$eV) [26]
@@
The BaF equilibrium configuration (CCSD, $m_{a}=1..10^{2}$eV) [25]
@@
Our results show that, despite the long-range nature of the property in question, its dependence on the geometry of the molecule is very similar to the short-range eN-SPS interaction. However, it should be stressed that previous studies of the axion-mediated interactions showed that electronic correlations can lead to significant corrections to the value of the enhancement parameters. This may increase the sensitivity of $W_{ax}$ to the vibrations of the molecule.
6 Acknowledgement

### v3

@@ V Results and discussion
This six terms correspond to a special anzats to represent a wave function in the core region. This choice of anzatz was made before in the works on the one-center restoration technique Titov et al. (2003) and used in the preceding work with same equivalent basis sets for Radium atom Zakharova and Petrov (2021). This is important for an accurate description of the short-range electron-nucleon interactions. However, in our case, because of the significant long-range contribution, this region contribution is negligible when $m_{a}\ll 1{\rm MeV}$. This is in sharp contrast to the eN-SPS interaction.
The equilibrium geometry was determined in Zakharova and Petrov (2021) to correspond to $R=4.4\,{\rm a.u.}$ The property was computed on a grid of values of $R$ and $\theta$. $R$ was taken to span values from $3.6\,{\rm a.u.}$ to $6.0\,{\rm a.u.}$ with step $0.2\,{\rm a.u.}$, where $\theta$ was taken to be $0^{\circ}$, $25^{\circ}$, $57^{\circ}$, $90^{\circ}$, $122^{\circ}$, $155^{\circ}$ and $180^{\circ}$. The resulting value for the equilibrium geometry equals $1.1407\cdot 10^{-5}\lambda_{e}^{-1}$ where $\lambda_{e}=\frac{\hbar}{m_{e}c}$ is the electron Compton length. Compare this with $3.36\cdot 10^{-5}\lambda_{e}^{-1}$ for YbOH obtained in Maison et al. (2021a). The dependence on the $\theta$ and $R$ is depicted on the Fig. 3 and Fig. 4. The values for several rovibrational states are given in Table 1. To check the convergence and estimate numerical error the computation was performed for $l_{max},j_{max},k_{max}=40,\ldots 43$, and $n_{max}=10,\ldots 13$.
@@
The YbOH equilibrium (FS-CCSD, $m_{a}=1..10$eV) Maison et al. (2021a)
@@
The BaF equilibrium (CCSD, $m_{a}=1..10^{2}$eV) Prosnyak and Skripnikov (2024)
@@
Our results show that, despite the long-range nature of the property in question, its dependence on the geometry of the molecule is very similar to the short-range eN-SPS interaction. However, it should be stressed that previous studies of the axion-mediated interactions showed that electronic correlations can lead to significant corrections to the value of the molecular parameters. This may increase the sensitivity of $W_{ax}$ to the vibrations of the molecule.
To test the validity of our computation we have also computed the values for different masses $m_{a}$ to compare the result with the prediction using the approximation (9) and the value for $W_{s}$ taken from Zakharova and Petrov (2021). As both properties shown very similar behavior we have restricted ourselves to the equilibrium configurations. The dependence is depicted on Fig. 5. For $m_{a}>5\cdot 10^{7}eV$ the deviation from predictions (9) are less than 10%. Note that just as in Prosnyak and Skripnikov (2024) we see a change of the sign of the property near $m_{a}\sim 10^{4}{\rm eV}$ which results in the blind spot for this mass range.
Finally on Fig. 6 we contrast the expected sensitivity of the RaOH experiment with the existing laboratory limits on the product of the axion-nucleon scalar and axion-electron pseudoscalar coupling constants O’Hare and Vitagliano (2020). Current experiments with diatomic molecules putting a limit $d_{e}\leq 1.3\cdot 10^{-30}\,e\cdot{\rm cm}$ for $\mathcal{E}=23\,{\rm GV}/{\rm cm}$ are able to sense energy differences $\Delta E=d_{e}\mathcal{E}\sim 0.7\cdot 10^{-5}Hz$ Roussy et al. (2023) which gives us the conservative estimate for the RaOH experiment constraint. It is expected Kozyryev and Hutzler (2017) that for the triatomic molecules this sensitivity may be improved by about two orders of magnitude which yields the optimistic constraint. This is still not enough to beat the astrophysical constraints coming from the red giants $g_{e,P}g_{N,S}<10^{-26}$, however this would make such molecular experiments best among laboratory experiments for the mass range $m_{a}>1\rm{eV}$.
Radium is a heavier atom than Ytterbium, which have resulted in higher values for sensitivities to such interactions as eEDM and eN-SPS for RaOH compared to YbOH molecules Zakharova and Petrov (2021); Zakharova (2022). One could expect a higher value for the considered axion-mediated effect in the RaOH molecule, but this is not a case. Hence, while our result shows that the RaOH molecule is less sensitive to this particular interaction, this result nevertheless is interesting. On the other hand, it is important to know such parameters for future experiment. Considering a different ratio with YbOH value of this parameter compared to the eEDM and eN-SPS parameters, performing the measurement with different molecules would allow to distinguish different P,T-odd effects from each others. While eEDM and eN-SPS are amplified for the RaOH compared to the YbOH by a similar factor, the axion-mediated interaction is significantly reduced and this could facilitate this separation.
Figure 5: The dependence of $W_{ax}$ on $m_{a}$
Figure 6: The existing laboratory constraints on the $g_{e,P}g_{N,S}$ and the projected RaOH constraints. The conservative constraint corresponds to the present level of energy split measurement achieved on diatomics, whereas the optimistic ones are given by the expected increase of precision on the triatomic molecules.
VI Acknowledgement